OpenStudy (anonymous):

a tank can provide water for 50 families for 5 days. how many days longer can the same tank provide water for 20 families?

OpenStudy (dan815):

Depends if the family lives

OpenStudy (dan815):

lol sorry!

OpenStudy (dan815):

50 families *5 days = 20 families * x days

OpenStudy (dan815):

x=50*5/20=25/2

OpenStudy (anonymous):

If a tank of water can provide 50 families with water for 5 days then that same tank could provide 20 families with 8 days of water. Think of this question as a ratio. 50:5 (50 families to 5 days) Since this ratio simplifies to 10:1 that says that for every 10 families taken away you gain an additional day that the said tank of water lasts. 50 Families use a tank of water that lasts 5 days40 Families use a tank of water that lasts 6 days30 Families use a tank of water that lasts 7 days20 Families use a tank of water that lasts 8 days.

OpenStudy (dan815):

-.-

OpenStudy (anonymous):

but the choices are 6, 6 1/2, 7 , 7 1/2?

OpenStudy (anonymous):

hey dan815 kindly explain your work to me .. thanks..

OpenStudy (dan815):

@ikram002p explain plz

OpenStudy (anonymous):

Cause they camels, they only drink once a week

OpenStudy (dan815):

lol

OpenStudy (dan815):

hmm those are the choices really?

OpenStudy (anonymous):

guyz, it is really hard answering this question kindly explain it to me guyz..

OpenStudy (anonymous):

yes..

OpenStudy (dan815):

its like this.. if 5 people eat 5 apples

OpenStudy (dan815):

then thats 1 apple for 1 person

OpenStudy (dan815):

if 2 people have to eat 5 apples then thats 5/2 apples per person

OpenStudy (anonymous):

yes.. i got you then?

OpenStudy (dan815):

they only need to eat 1 apple a day, so it will last them 5/2 times longer

OpenStudy (dan815):

because an apple a day keeps the doctor away

OpenStudy (anonymous):

But nowadays to get the Nutrients in an apple back in 1850, you will have to eat 5 apples.

OpenStudy (anonymous):

that one i dont really understand..

OpenStudy (ikram002p):

is it linear relation ?

OpenStudy (anonymous):

i also dont know whta topic is this..

OpenStudy (ikram002p):

the tank is constant variable , so just make a linear relation btw families and days